7. A cylinder containing the air comprises the system. Cycle is completed as

follows :

(i) 82000 N-m of work is done by the piston on the air during compression stroke and

45 kJ of heat are rejected to the surroundings.

(ii) During expansion stroke 100000 N-m of work is done by the air on the piston.

Calculate the quantity of heat added to the system.


1
Expert's answer
2021-09-07T16:31:21-0400


Compression stroke. Process 1-2 :

Work done by the piston on the air, "W_{1–2} = – 82000 \;Nm\; (= – 82 \;kJ)"

Heat rejected to the system,

"Q_{1–2} = – 45 \;kJ \\

Q_{1–2} = (U_2 - U_1 ) + W \\

– 45 = (U_2 - U_1 ) + (– 82) \\

(U_2-U_1) = 37 \;kJ"

Work done by air on the piston,

"W_{2–1} = 100000 \;Nm \;(= 100 \;kJ) \\

Q_{2–1} = (U_1 - U_2 ) + W \\

= – 37 + 100\; kJ = 63\; kJ"

Hence, quantity of heat added to the system = 63 kJ.